In 1993, the number of Americans dying from cancer was 6 times the number that died from accidents. If the number of deaths from these two causes totaled 630,000 how many americans died from each cause?
Answer by checkley77(12569)

X+6X=630,000
7X=630,000
X=630,000/7
X=90,000 DIED IN ACCIDENTS.
6*90,000=540,000 DIED OF CANCER.
PROOF:
90,000+540,000=630,000
630,000=630,000